# Relative Distance Guided Dynamic Partition Learning for Scale-Invariant UAV-View Geo-Localization

### v2

@@ III Methods
Task definition. The geo-localization dataset contains images pairs captured from different platforms, defined as $\{x_{D},x_{S}\}$, where subscripts $D$ and $S$ indicate drone- and satellite-view, respectively. We denote the label as $y\in[1,C]$, where $C$ indicates the number of categories. For instance, the label of University-1652 [10] dataset is $y\in[1,701]$. For UAV-view geo-localization, we intend to learn a mapping function that projects images from various platforms to one shared semantic space. Image pairs with the same geo-tag have a consistent spatial distribution (i.e., close to each other), otherwise, the image pairs are far apart from each other.
Figure 4: Overview of DGDPL framework, including three phase: feature extraction, altitude guided dynamic partition learning and classification supervision. In the testing phase, part-level image representation is extracted before classification layer in classifier module, and measures similarity by Euclidean distance.
@@ III-A Feature Extraction
Both drone-view and satellite-view images fall within the category of aerial-view imagery, exhibiting comparable feature domains. Consequently, We employ one feature extractor to process cross-view images. As mentioned before, our DGDPL belongs to hard-partition learning and is compatible with various backbone networks, such as ResNet [80] and ViT [79]. Following previous works [24, 55], we adopt ResNet-50 for illustration. Given an image $x_{i}\in\mathbb{R}^{3\times 512\times 512}$, after the backbone network and an upsampling layer we can acquire the corresponding high-level feature map $f_{i}\in\mathbb{R}^{2048\times 32\lambda\times 32\lambda}$. The process can be formulated as:
$f_{i}=\mathcal{U}_{\lambda}(\mathcal{F}_{backbone}(x_{i})),i\in\{D,S\}$ (1)
@@
where $\mathcal{F}_{backbone}$ stands for the process of feature extraction, and $\mathcal{U}_{\lambda}$ stands for the upsampling process with scale factor $\lambda$. We emphasize that the degree of variation in partition size can be refined by upsampling feature map $f_{i}$.
III-B Distance Guided Dynamic Partition Learning
@@
To capture aligned contextual information even in scale-inconsistent scenes, we propose a distance guided dynamic partition learning, consisting of a square partition strategy, a distance-guided adjustment strategy, and a saliency-guided refinement strategy.
@@
Square partition strategy (SPS). We propose a simple yet effective partition strategy to generate fine-grained features. As shown in Case(II) of Fig. 2, our square partition strategy splits the input feature into $N_{SPS}$ parts following square templates. Concretely, after SPS processing, the high-level feature map $f_{i}\in\mathbb{R}^{2048\times 32\lambda\times 32\lambda}$ is segmented into $N_{SPS}$ part-level feature $f_{i}^{n_{SPS}}\in\mathbb{R}^{2048\times\frac{32\lambda n_{SPS}}{N_{SPS}}\times\frac{32\lambda n_{SPS}}{N_{SPS}}}(n_{SPS}\in[1,N_{SPS}])$, where $n_{SPS}$ denotes the $n$-th part. The overall SPS processing can be formulated as:
$\Omega(f_{i}^{n_{SPS}})=\{f_{i}^{n_{SPS}}|\mathcal{F_{SPS}}(f_{i};N_{SPS})\}$ (2)
where $\mathcal{F}_{SPS}(\cdot)$ stands for the process of SPS, and $\Omega(f_{i}^{n_{SPS}})$ denotes the set of all fine-grained features.
@@
Despite the similarities with LPN, by comparing Fig. 3 (b) and (c) we state that SPS has two distinct advantages: 1) SPS mines local features while maintaining global information, enhancing robustness against scale changes; 2) the solid partition is readily obtainable and its size can be easily adjusted.
@@
Distance-guided adjustment strategy (DGAS). Existing hard-partition methods utilize uniform templates for processing multi-view images. However, in scenarios where drone-view scale undergo significant variations, these methods exhibit notable performance degradation. The primary reason is the misalignment of semantic content within partition pairs. To solve this limitation, we propose an innovative distance-guided adjustment strategy, i.e., dynamically adjusting partition size in response to drone-view scale changes, thus boosting content consistency between partition pairs.
Specifically, both the pose of UAVs and the resolution of satellite maps belong to additional attributes which can be obtained in the GPS-denied environment. Based on these attributes, the distance between drone/satellite-views and the ground can be estimated, linearly related to the image scale. In this paper, we assume that the distance between drone/satellite-views to the ground is known. Then, we introduce a scale factor $\beta$ to quantify the relative distance between cross-view images:
$\beta=\frac{H_{D}-H_{S}}{H_{S}}$ (3)
where $H_{D}$ and $H_{S}$ denote the distance from drone-view and satellite-view to the ground, respectively. Note that the satellite-view distance $H_{S}$ is constant, which is the regular setting for UVGL.
Since modifying partitions in high-dimensional space instead of RGB domain, we further construct the adjustment factor $\theta$ based on $\beta$:
$\theta=int(\beta\cdot\lambda\cdot\alpha)$ (4)
where $int(\cdot)$ stands for the round operation, $\lambda$ denotes the upsampling factor mentioned in Eq.(1), and $\alpha$ denotes the hyper-parameter utilized to regulate the degree of partition change. The sign of $\theta$ is determined by $\beta$, where ($\theta$$>$0) denotes indenting partitions, whereas ($\theta$$<$0) denotes expanding partitions.
@@
Then, we inject $\theta$ into SPS to adjust partitions of drone-views, which can be formulated as:
$\Omega(f_{D}^{n_{SPS}})=\{f_{D}^{n_{SPS}}|\mathcal{F_{SPS}^{D}}(f_{D};N_{SPS};\theta)\}$ (5)
@@
where $\mathcal{F_{SPS}^{D}}(\cdot)$ denotes the process of SPS specific for drone-views. For ease of understanding, we depict the $\mathcal{F_{SPS}^{D}}$ process in two cases (i.e., $\theta$$>$0 and $\theta$$<$0), as shown in Fig. 2 (Case I and III). We employ $\theta$ to adjust the size of partitions, and the shape of each partition pair is expressed as follows:
@@
$\left\{\begin{array}[]{l}f_{D}^{n_{SPS}}\in\mathbb{R}^{2048\times(\frac{32\lambda n_{SPS}}{N_{SPS}}-2\theta)\times(\frac{32\lambda n_{SPS}}{N_{SPS}}-2\theta)}\\ f_{S}^{n_{SPS}}\in\mathbb{R}^{2048\times(\frac{32\lambda n_{SPS}}{N_{SPS}})\times(\frac{32\lambda n_{SPS}}{N_{SPS}})}\\ \end{array}\right.$ (6)
Note that the global feature $f_{D}^{N_{SPS}}$ is invariant.
In practice, the size of partitions must be positive. In other words, when $\theta$$>$0, $\theta$ must adhere to the following constraint:
$max(|\theta|)\leq\frac{32\lambda}{2N_{SPS}}-1$ (7)
Based on Eq. (3), (4) and (7), we can derive the range of $\alpha$:
$0\leq\alpha\leq\frac{16\lambda-N_{SPS}}{N_{SPS}}\cdot\frac{1}{\lambda}\cdot\frac{H_{S}}{max(H_{D})-H_{S}}$ (8)
where $max(H_{D})$ denotes the maximum distance of drone-views. When $\alpha$$<$0, the extreme case is that all partitions expand equal to the global feature, thereby necessitating no constraint on the value range of $\alpha$. Hence it suffices to constrain $\alpha$ based on Eq. (8). Note that we omit the effect of lens focal length on the drone-view scale and default to a uniform lens setting for drone-view images. Therefore, $\alpha$ can be selected experimentally.
@@
Saliency-guided refinement strategy (SGRS). Each solid partition inherently contains both target and environment information. To further refine part-level features, we design a saliency-guided refinement strategy. The overall process are depicted in Fig. 4, and can be expressed as:
$(g_{i}^{n_{SPS}},s_{i}^{n_{SPS}},b_{i}^{n_{SPS}})=\mathcal{F}_{SGRS}(f_{i}^{n_{SPS}})$ (9)
@@
where $\mathcal{F}_{SGRS}(\cdot)$ denotes the process of SGRS, and $g_{i}^{n_{SPS}}$, $s_{i}^{n_{SPS}}$, $b_{i}^{n_{SPS}}$ are denote the output global feature, salient feature and background feature, respectively.
@@
Specifically, assuming the input partition feature is $f_{i}^{n_{SPS}}\in\mathbb{R}^{2048\times h\times w}$, we first generate the heatmap $heat_{i}^{n_{SPS}}\in\mathbb{R}^{1\times h\times w}$ to reflect importance regions of the input. Considering that targets are usually centered, we introduce an auxiliary coordinate map ($CM$) to direct the heatmap’s focus towards the central region. The $CM$ indicates inverse pixel-to-center distance, ranging from 1 (center) to 0 (farthest). To be consistent with the partition shape, we adopt Chebyshev distance to generate $CM$:
$CM(p,q)=max(\frac{h}{2},\frac{w}{2})-max(|p-\frac{h}{2}|,|q-\frac{w}{2}|)$ (10)
@@
where $(p,q)$ denotes the coordinates of pixels. Note that other distance metrics, e.g, Euclidean distance and Manhattan distance, will be compared in the experiments. The generation of $heat_{i}^{n_{SPS}}$ can be formulated as:
$heat_{i}^{n_{SPS}}=(\mathcal{N}(\mathcal{M}(f_{i}^{n_{SPS}}))+CM_{i}^{n_{SPS}})/2$ (11)
@@
where $\mathcal{N}(\cdot)$ denotes the normalization operation, and $\mathcal{M}(\cdot)$ denotes the average pooling operation. Next, we set a threshold $\delta$ to split the heatmap, resulting in a binary mask $Mask_{i}^{n_{SPS}}$:
$Mask_{i}^{n_{SPS}}(p,q)=\left\{\begin{array}[]{ll}1,&heat_{i}^{n_{SPS}}(p,q)\geq\delta\\ 0,&heat_{i}^{n_{SPS}}(p,q)<\delta\end{array}\right.$ (12)
@@
Finally, in reference to FSRA [19], we leverage the value of $Mask_{i}^{n_{SPS}}$ to binary the input feature $f_{i}^{n_{SPS}}$, which can be expressed as follows:
$\left\{\begin{array}[]{l}s_{i}^{n_{SPS}}=\mathcal{M}(\{f_{i}^{n_{SPS}}(p,q)|Mask_{i}^{n_{SPS}}(p,q)=1\})\\ b_{i}^{n_{SPS}}=\mathcal{M}(\{f_{i}^{n_{SPS}}(p,q)|Mask_{i}^{n_{SPS}}(p,q)=0\})\\ g_{i}^{n_{SPS}}=\mathcal{M}(f_{i}^{n_{SPS}})\end{array}\right.$ (13)
@@
where $s_{i}^{n_{SPS}}\in\mathbb{R}^{2048\times 1\times 1}$ and $b_{i}^{n_{SPS}}\in\mathbb{R}^{2048\times 1\times 1}$ are denote the salient feature and background feature. We also utilize $g_{i}^{n_{SPS}}\in\mathbb{R}^{2048\times 1\times 1}$ to preserve the input information. It is worth mentioning that SGRS is implemented for each part-level features of both drone-view and satellite-view, resulting in $3N_{SPS}$ feature vectors.
@@ III-C Classification Supervision
We employ a classifier module to map feature vectors of all sources into one shared space, As shown in Fig. 4, the classifier module includes a fully connected layer (FC), a batch normalization layer (BN), a dropout layer (Dropout), and a classification layer (Cls). The process can be expressed as:
$z_{i,j}^{n_{SPS}}=\mathcal{F}_{classifier}(j_{i}^{n_{SPS}})$ (14)
where $j\in\{g,s,b\}$ denotes the feature vector of different categories generated by Eq. (13).
Subsequently, we adopt a cross-entropy loss function to optimize our framework, with the objective of minimizing the distance between features of the same geo-tag:
$\mathcal{L}_{CE}=\sum_{i,j,n_{SPS}}-log(\frac{exp(z_{i,j}^{n_{SPS}}(y))}{\sum_{c=1}^{C}(z_{i,j}^{n_{SPS}}(c))})$ (15)
where $z_{i,j}^{n_{SPS}}(y)$ denotes the logit score of the ground-truth geo-tag $y$, and $C$ denotes the number of geo-tag categories. It is noteworthy that the losses on the image of different part-level features are calculated independently.
IV Experiments
To evaluate the effectiveness of our DGDPL framework for UAV-view geo-localization, we conduct a series of experiments on University-1652 [10]. These experiments assess the performance of our method in terms of retrieval accuracy, robustness against scale variations, and the impact of both various components and hyper-parameters.

### v3

@@ III Methodology
Task definition. The geo-localization dataset contains image pairs captured from different platforms, defined as $\{x_{D},x_{S}\}$, where subscripts $D$ and $S$ indicate drone and satellite-platform, respectively. We denote the label as $y\in[1,C]$, where $C$ indicates the number of categories. For instance, the label of the University-1652 [8] dataset is $y\in[1,701]$. For UAV-view geo-localization, we intend to learn a mapping function that projects images from various platforms to one shared semantic space. Image pairs with same labels are close to each other, and vice versa are far away. Therefore, given the drone-view images, user could quickly search the corresponding images from other platform during the inference time. According to the meta-data of other platform, we localize the drone even in the GNSS-denied areas.
Overview. As depicted in Fig. 4, SaLPN consists of three stages: feature extraction, scale-adaptive partition learning and classification supervision. In the feature extraction stage, a backbone network maps drone- and satellite-view images to high-level feature maps. Then, we introduce scale-adaptive partition learning to generate fine-grained features with consistent content. We further propose a saliency-guided refinement strategy to classify each part-level feature, thus obtaining finer discriminative representations. Finally, we adopt a classifier to predict the category of features obtained from different views, followed by using cross-entropy loss to optimize the model.
[…]
@@ III-A Feature Extraction
Both drone-view and satellite-view images fall within the category of aerial-view images, exhibiting comparable feature domains. Consequently, we employ one feature extractor to process cross-view images. As mentioned before, SaLPN belongs to hard-partition learning and is compatible with various backbone networks, such as ResNet [73] and ViT [72]. Following previous works [25, 53], we adopt ResNet-50 for illustration. Given an image $x_{i}\in\mathbb{R}^{3\times 512\times 512}$, after the backbone network and an upsampling layer, we can acquire the corresponding high-level feature map $f_{i}\in\mathbb{R}^{2048\times 128\times 128}$. The process can be formulated as:
$\left\{\begin{array}[]{c}f_{D}=\mathcal{U}_{\times 4}(\mathcal{F}_{backbone}(x_{D})),\\ f_{S}=\mathcal{U}_{\times 4}(\mathcal{F}_{backbone}(x_{S})),\end{array}\right.$ (1)
@@
where $f_{D}$ and $f_{S}$ are the high-level features of the drone-view and satellite-view, respectively; $\mathcal{F}_{backbone}$ stands for the process of feature extraction; $\mathcal{U}_{\times 4}$ stands for a ×$\times 4$ up-sampling layer. We emphasize that the degree of variation in partition size can be refined by upsampling feature maps.
III-B Scale-adaptive Partition Learning
@@
To capture aligned contextual information even in scale-inconsistent scenes, we propose a scale-adaptive partition learning, consisting of a square partition strategy, a height-aware adjustment strategy, and a saliency-guided refinement strategy.
@@
Square partition strategy (SPS). We propose a simple yet effective partition strategy to generate fine-grained features. As shown in Case(II) of Fig. 2, SPS splits the input feature into $N$ parts following square templates. Concretely, after SPS processing, the high-level feature map $f_{i}\in\mathbb{R}^{2048\times 128\times 128}$ is segmented into $N$ part-level feature $f_{i}^{n}\in\mathbb{R}^{2048\times\frac{128n}{N}\times\frac{128n}{N}}(n\in[1,N])$, where $n$ denotes the $n$-th part. The conventional SPS processing can be formulated as:
$\Omega(f_{i}^{n})=\{f_{i}^{n}|\mathcal{F_{SPS}}(f_{i};N)\},$ (2)
@@
where $\mathcal{F}_{SPS}(\cdot)$ stands for the process of SPS, and $\Omega(f_{i}^{n})$ denotes the set of all part-level features. Despite the similarities with LPN, comparing Fig. 3 (b) and (c) shows that SPS mines local features while maintaining global information, enhancing robustness against scale changes.
@@
Height-aware adjustment strategy (HAAS). Existing hard-partition methods adopt uniform templates for processing multi-view images. However, in scenarios where drone-view scale undergo significant variations, these methods exhibit notable performance degradation. The primary reason is the misalignment of semantic content within partition pairs. To alleviate this limitation, we propose an innovative height-aware adjustment strategy, i.e., dynamically adjusting partition size in response to drone-view scale changes, thus boosting content consistency between partition pairs.
First, we assume that the shooting height of drone-views is acquirable, while the height of satellite-view images can be estimated by comparing the area per unit pixel with a fixed height drone-view image. To quantify the change in partition size, we further construct a scale factor $\theta$ based on the relative height between cross-view images:
$\theta=\text{round}(\frac{H_{D}-H_{S}}{H_{S}}\cdot\alpha),$ (3)
where $H_{D}$ and $H_{S}$ denote the height from drone-view and satellite-view to the ground, respectively; $\text{round}(\cdot)$ stands for the rounding operation, and partition adjustment factor $\alpha$ denotes a hyper-parameter utilized to regulate the degree of partition change. ($\theta$$>$0) indicates that the partition needs to be decreased, and ($\theta$$<$0) vice versa. Note that the height of satellite-view $H_{S}$ is constant, which is the regular setting for UVGL.
@@
Then, different from the traditional SPS in Eq. 2, we further take $\theta$ into consideration, which adjusts partitions of drone-views, which can be formulated as:
$\Omega(f_{D}^{n})=\{f_{D}^{n}|\mathcal{F_{HAAS}^{D}}(f_{D};N;\theta)\},$ (4)
@@
where $\mathcal{F_{HAAS}^{D}}(\cdot)$ denotes the process of HAAS specific for drone-views. For ease of understanding, we depict the $\mathcal{F_{HAAS}^{D}}$ process in two cases (i.e., $\theta$$>$0 and $\theta$$<$0), as shown in Fig. 2 (Case I and III). We employ $\theta$ to adjust the size of partitions, and the shape of each partition pair is expressed as follows:
@@
$\left\{\begin{array}[]{l}f_{D}^{n}\in\mathbb{R}^{2048\times(\frac{128n}{N}-2\theta)\times(\frac{128n}{N}-2\theta)},\\ f_{S}^{n}\in\mathbb{R}^{2048\times(\frac{128n}{N})\times(\frac{128n}{N})}.\end{array}\right.$ (5)
In practice, there are two extreme cases that should be emphasized: 1) when $\theta$$>$0, the size of the smallest partition must be positive; 2) while $\theta$$<$0, the minimum partition block expands into global features, i.e., all partitions are equal to the global feature. This means that $\alpha$ in Eq. 3 has a limited range when the height range of the dataset is known. Next, we will derive the range of values of $\alpha$.
Specifically, when $\theta$$>$0 (i.e., $H_{D}$$>$$H_{S}$), we have $1\leq\frac{128}{2N}-\theta$ to avoid non-positive partitions. Based on Eq. 3, we can derive the range of the hyperparameter $\alpha$:
$0\leq\alpha\leq\frac{64-N}{N}\cdot\frac{H_{S}}{H_{D}^{max}-H_{S}},$ (6)
where $H_{D}^{max}$ denotes the maximum height of drone-views.
Conversely, when $\theta$$<$0 (i.e., $H_{D}$$<$$H_{S}$), we have $\frac{128}{2N}-\theta\leq\frac{128}{2}$. Based on Eq. 3, we also can derive the range of the hyperparameter $\alpha$:
$0\leq\alpha\leq\frac{64(N-1)}{N}\cdot\frac{H_{S}}{H_{S}-H_{D}^{min}},$ (7)
where $H_{D}^{min}$ denotes the minimum height of drone-views.
If all same-height image pairs have $H_{S}=H_{D}$, we could derive $\theta=0$, which is independent from the selection of $\alpha$. Therefore, for a dataset with a clear height range, the value of $\alpha$ needs to satisfy both Eq. 6 and 7.
@@
Saliency-guided refinement strategy (SGRS). Each solid partition inherently contains both target and environment information. To further refine part-level features, we design a saliency-guided refinement strategy. The overall process is depicted in Fig. 4, and can be expressed as:
$(g_{i}^{n},s_{i}^{n},b_{i}^{n})=\mathcal{F}_{SGRS}^{n}(f_{i}^{n}),$ (8)
@@
where $\mathcal{F}_{SGRS}^{n}(\cdot)$ denotes the process of SGRS, and $g_{i}^{n}$, $s_{i}^{n}$, $b_{i}^{n}$ are denote the output global feature, salient feature and background feature, respectively.
@@
Specifically, assuming the input partition feature is $f_{i}^{n}\in\mathbb{R}^{2048\times h\times w}$, we first generate the heatmap $heat_{i}^{n}\in\mathbb{R}^{1\times h\times w}$ to reflect importance regions of the input. Considering that targets are usually centered, we introduce an auxiliary coordinate map ($CM$) to direct the heatmap’s focus towards the central region. $CM$ indicates inverse pixel-to-center distance, ranging from 1 (center) to 0 (farthest). To be consistent with the partition shape, we adopt Chebyshev distance to generate $CM$:
$CM(u,v)=max(\frac{h}{2},\frac{w}{2})-max(|u-\frac{h}{2}|,|v-\frac{w}{2}|),$ (9)
@@
where $(u,v)$ denotes the coordinates of pixels. Note that other distance metrics, e.g, Euclidean distance and Manhattan distance, will be compared in the experiments. The generation of $heat_{i}^{n}$ can be formulated as:
$heat_{i}^{n}(u,v)=(Norm(Avg(f_{i}^{n}(u,v)))+CM(u,v))/2,$ (10)
@@
where $Norm(\cdot)$ denotes the normalization operation, and $Avg(\cdot)$ denotes the average pooling operation. Next, we set a threshold $\delta$ to split the heatmap, resulting in a binary mask $Mask_{i}^{n}$:
$Mask_{i}^{n}(u,v)=\left\{\begin{array}[]{ll}1,&heat_{i}^{n}(u,v)\geq\delta,\\ 0,&heat_{i}^{n}(u,v)<\delta.\end{array}\right.$ (11)
@@
Finally, in reference to FSRA [20], we leverage the value of $Mask_{i}^{n}$ to split the input feature $f_{i}^{n}$, which can be expressed as follows:
$\left\{\begin{array}[]{l}s_{i}^{n}=Avg(f_{i}^{n}(u,v)\odot Mask_{i}^{n}(u,v)),\\ b_{i}^{n}=Avg(f_{i}^{n}(u,v)\odot(1-Mask_{i}^{n}(u,v))),\\ g_{i}^{n}=Avg(f_{i}^{n}),\end{array}\right.$ (12)
@@
where $\odot$ denotes the pixel-level multiplication, $s_{i}^{n}\in\mathbb{R}^{2048\times 1\times 1}$ and $b_{i}^{n}\in\mathbb{R}^{2048\times 1\times 1}$ denote the salient feature and background feature, respectively. We also extract $g_{i}^{n}\in\mathbb{R}^{2048\times 1\times 1}$ to preserve the input information. It is worth mentioning that SGRS is implemented for each part-level features of both drone-view and satellite-view, resulting in $3N$ feature vectors.
@@ III-C Classification Supervision
We employ 3N classifier modules to map feature vectors of all sources into three corresponding shared space. As shown in Fig. 4, the classifier module includes a fully connected layer (FC), a batch normalization layer (BN), a dropout layer (Dropout), and a classification layer (Cls). The process can be expressed as:
$p_{g}^{i,n}=\mathcal{F}_{g}^{n}(g_{i}^{n}),\quad p_{s}^{i,n}=\mathcal{F}_{s}^{n}(s_{i}^{n}),\quad p_{b}^{i,n}=\mathcal{F}_{b}^{n}(b_{i}^{n}),$ (13)
where $\mathcal{F}_{g}^{n},\mathcal{F}_{s}^{n},\mathcal{F}_{b}^{n}$ are three independent classifier modules, respectively. It is worth noting that the classifiers adopted for different granularity feature vectors are also independent. Subsequently, we adopt a cross-entropy loss function to optimize our framework, with the objective of minimizing the distance between features of the same geo-tag:
$\mathcal{L}_{CE}=\sum_{i,n}-log(p_{g}^{i,n}(y))-log(p_{s}^{i,n}(y))-log(p_{b}^{i,n}(y)),$ (14)
where $p_{g}(y)^{i,n}$ denotes the logit score of the ground-truth geo-tag $y$ based on the global feature $g^{n}_{i}$. Similarly, $p_{s}^{i,n}(y)$ and $p_{b}^{i,n}(y)$ are the predicted probability of the ground-truth location based on the salient and background feature, respectively.
IV Experiment
To evaluate the effectiveness of SaLPN for UAV-view geo-localization, we conduct comparative experiments on the University-1652 and SUES-200 datasets, and ablation experiments on the University-1652 dataset. These experiments assess the performance of our method in terms of retrieval accuracy, robustness against scale variations, and the impact of both various components and hyper-parameters.
Figure 5: The data augmentation strategy to extend the shooting height range of drone-views from the visual perspective. (a) Making the shooting height seem higher: given a drone image, mirror a strip $\Delta$P pixels wide around the input image and resize it to initial resolution; (b) Making the shooting height seem lower: given a drone image, crop out a square ring $\Delta$P pixels wide around the input image and resize it to initial resolution.